OpenStudy (anonymous):

Which expression is equivalent to sec2xcot2x?

OpenStudy (anonymous):

\[\frac{ 1 }{ \cos ^{2}x }\] \[\frac{ \cos ^{2}x }{ \sin ^{2}x }\]

OpenStudy (anonymous):

multiply them and you get \[\frac{ 1 }{ \sin ^{2}x }=\csc ^{2}x\]